# Adversarial Attacks for Drift Detection

### v1

@@ A.2 Proofs of the statements in Table 1
$\ker\mathcal{W}=\left\{f\in L^{1}_{\text{loc}}(\mathcal{T})\>\left|\>\lambda(W_{2})\int_{W_{1}}f(t)\textnormal{d}t=\lambda(W_{1})\int_{W_{2}}f(t)\textnormal{d}t\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}.$
@@
For $\mathcal{W}=\{([0,t],[t,t+l])\mid a\leq t\}$ with $a,l>0$ and $f\in L^{1}_{\text{loc}}(\mathbb{R}_{\geq 0})$ the following are equivalent
@@
$C$ $:=\frac{l}{a}\int_{0}^{a}f(t)\textnormal{d}t$ and
$F_{t}(n)$ $:=\int_{0}^{a+t+ln}f(t)\textnormal{d}t.$
@@
For $\mathcal{W}=\{([t-l,t],[t,t+l])\mid t\in\mathbb{R}\}$ with $l>0$ and $f\in L^{1}_{\text{loc}}(\mathbb{R})$ the following are equivalent
@@
we conclude that $c_{2}$ is continues, too.
@@
we see the quotient of the form $P_{h}(t)+tQ(t)+o(h)$ for all $h>0$ with $P$ and $Q$ $l$-periodic since $c_{1}$ and $c_{2}$ are $l$-periodic. On the one hand, have $(F_{t+h}(0)-F_{t-h}(0))/(2h)\to f(t)$ as $h\to 0$ by Lebesgue’s differentiation theorem so $f(t)=p(t)+q(t)t$ with $p$ and $q$ $l$-periodic.
@@
2. $\Leftrightarrow$ 3. if 3. holds or – according to the previous claim – if 2. holds, we have $f(t)=p(t)+tq(t)$ with $p,q\in L^{1}_{\text{loc}}(\mathbb{R})$ $l$-periodic. Therefore, it remains to show that $3.\Leftrightarrow\int_{0}^{l}q(x)\textnormal{d}x=0$. It holds
@@
Let $\mathcal{W}=\{([0,t],[t,t+l])\mid t\geq a\}$ with $a,l>0$ and $f\in L^{1}_{\text{loc}}(\mathbb{R}_{\geq 0})$. The following are equivalent:
@@
    $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+(t+a)q(t)$ for all $t>0$ with $p$ and $q$ $l$-periodic, $p(t)=\int_{0}^{t}q(x)\textnormal{d}x+C$, $\int_{0}^{l}q(x)\textnormal{d}x=0$, and $\int_{0}^{a}g(x)\textnormal{d}x=0$.
In particular, there are boundary effects, i.e., $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}\subsetneq\ker\mathcal{W}$, and $\textnormal{Adv}_{0}(\mathcal{W})$ are exactly the functions that are arbitrary before $a$ and then constant with the same mean as before, i.e., $a^{-1}\int_{0}^{a}f(x)\textnormal{d}x=f(t)$ for all $t>a$.
@@
2. $\Rightarrow$ Properties of 3. except $\int_{0}^{a}g(x)\textnormal{d}x=0$ define $F_{t}(n)=\int_{0}^{a+t+ln}f(x)\textnormal{d}x$ then we have
$(a+t+l(n-1))^{-1}F_{t}(n-1)=l^{-1}(F_{t}(n)-F_{t}(n-1))$
which allows the recursive definition for each $t\in\mathbb{R}_{\geq 0}$
$F_{t}(n)$ $=\frac{l+l(n-1)+a+t}{l(n-1)+a+t}F_{t}(n-1)$
$=\underbrace{\frac{ln+a+t}{l(n-1)+a+t}\frac{l(n-1)+a+t}{l(n-2)+a+t}}_{=\frac{ln+a+t}{l(n-2)+a+t}}F_{t}(n-2)$
$=\cdots=\frac{ln+a+t}{a+t}F_{t}(0)$
$=(t+ln+a)\underbrace{\frac{1}{a+t}F_{t}(0)}_{=:c(t)}$
@@
that since $t\mapsto F_{t}(n)$ is continuous that $c(t)$ is continuous.
@@
$\frac{F_{t+h}(0)-F_{t-h}(0)}{2h}$ $=\left(t-h+a\right)\frac{c(t+h)-c(t-h)}{2h}+c(t+h).$
Thus, by Lebesgue’s differentiation theorem, we then conclude that $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+(t+a)q(t)$ with $p,q$ $l$-periodic, and $p(t)=C+\int_{0}^{t}q(x)\textnormal{d}x$ for all $t>0$ and $g$ arbitrary. In particular, as $p$ is $l$-periodic we have $0=p(l)-p(0)=\int_{0}^{l}q(x)\textnormal{d}x$.
@@
As before we conclude that if $|f|\leq 1$ we have $q(t)=0$ and hence $p(t)=C$ is constant for all $t\geq a$ and $a^{-1}\int_{0}^{a}p(x)\textnormal{d}x=l^{-1}\int_{a}^{a+l}p(x)\textnormal{d}x=C$. ∎

### v2

@@ A.2 Proofs of the statements in Table 1
$\ker\mathcal{W}=\left\{f\in L^{1}_{\textnormal{loc}}(\mathcal{T})\>\left|\>\lambda(W_{2})\int_{W_{1}}f(t)\textnormal{d}t=\lambda(W_{1})\int_{W_{2}}f(t)\textnormal{d}t\forall(W_{1},W_{2})\in\mathcal{W}\right.\right\}.$
@@
For $\mathcal{W}=\{([0,a],[t,t+l])\mid a\leq t\}$ with $a,l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R}_{\geq 0})$ the following are equivalent
@@
$C$ $:=\frac{l}{a}\int_{0}^{a}f(x)\textnormal{d}x$ and
$F_{t}(n)$ $:=\int_{0}^{a+t+ln}f(x)\textnormal{d}x.$
@@
To prove the next statement we need the following simple lemma:
Lemma 1.
Let $a_{n},b_{n}\subset\mathbb{R}$ and $c_{1},c_{2}\in\mathbb{R}$ with $c_{1}\neq c_{2}$. The following are equivalent
  1. 1.
    The limits $a_{n}$ and $b_{n}$ as $n\to\infty$ exist
  2. 2.
    The limits $a_{n}+c_{1}b_{n}$ and $a_{n}+c_{2}b_{n}$ exist as $n\to\infty$
Proof.
1. $\Rightarrow$ 2. is just linearity.
2. $\Rightarrow$ 1. follows by considering $(a_{n}-c_{1}b_{n})-(a_{n}-c_{2}b_{n})=(c_{1}-c_{2})b_{n}$. As the limit $n\to\infty$ on the left hand side exists so does the limit on the right hand side which by linearity implies that the limit of $b_{n}$ exists. Thus, by linearity, the limit of $a_{n}$ exists, too. ∎
@@
For $\mathcal{W}=\{([t-l,t],[t,t+l])\mid t\in\mathbb{R}\}$ with $l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R})$ the following are equivalent
@@
we conclude that $c_{2}$ and hence $c_{1}$ are continues, too.
@@
$=\quad\underbrace{\frac{c_{1}(t+h)-c_{1}(t-h)}{2h}+c_{2}(t+h)}_{=:P(t)}$
$\quad+t\cdot\underbrace{\frac{c_{2}(t+h)-c_{2}(t-h)}{2h}}_{=:Q(t)}$
$\quad+\underbrace{\frac{c_{2}(t+h)-c_{2}(t-h)}{2}}_{\in o(h)}$
we see the quotient is of the form $P(t)+tQ(t)+o(h)$ for all $h>0$ with $P$ and $Q$ $l$-periodic since $c_{1}$ and $c_{2}$ are $l$-periodic. On the one hand, have $(F_{t+h}(0)-F_{t-h}(0))/(2h)\to f(t)$ as $h\to 0$ by Lebesgue’s differentiation theorem. As $P$ and $Q$ are $l$-periodic we obtain the same sequences for $a_{n}:=P(t+h_{n})=P(t+h_{n}+l)$ and $b_{n}:=Q(t+h_{n})=Q(t+h_{n}+l)$ so by Lemma 1 with $c_{1}=t$ and $c_{2}=t+l$ we see that the (point-wise) limits exist and thus obtain $l$-periodic $L^{1}_{\textnormal{loc}}$-function $p$ and $q$ with $f(t)=p(t)+q(t)t$.
@@
2. $\Leftrightarrow$ 3. if 3. holds or – according to the previous claim – if 2. holds, we have $f(t)=p(t)+tq(t)$ with $p,q\in L^{1}_{\textnormal{loc}}(\mathbb{R})$ $l$-periodic. Therefore, it remains to show that $3.\Leftrightarrow\int_{0}^{l}q(x)\textnormal{d}x=0$. It holds
@@
Let $\mathcal{W}=\{([0,t],[t,t+l])\mid t\geq a\}$ with $a,l>0$ and $f\in L^{1}_{\textnormal{loc}}(\mathbb{R}_{\geq 0})$. The following are equivalent:
@@
    $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+tq(t)$ for all $t>0$ with $p$ and $q$ $l$-periodic, $p(t)=\int_{0}^{t}q(x)\textnormal{d}x+C$, $\int_{0}^{l}q(x)\textnormal{d}x=0$, and $\int_{0}^{a}g(x)\textnormal{d}x=0$.
In particular, there are boundary effects, i.e., $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}\subsetneq\ker\mathcal{W}$, and $\textnormal{Adv}_{0}(\mathcal{W})\mathbb{R}$ are exactly the functions that are arbitrary before $a$ and then constant with the same mean as before, i.e., $a^{-1}\int_{0}^{a}f(x)\textnormal{d}x=f(t)$ for all $t>a$.
@@
2. $\Rightarrow$ Properties of 3. except $\int_{0}^{a}g(x)\textnormal{d}x=0$ define $F_{t}(n)=\int_{0}^{t+ln}f(x)\textnormal{d}x$ then we have
$(t+l(n-1))^{-1}F_{t}(n-1)=l^{-1}(F_{t}(n)-F_{t}(n-1))$
for all $t\geq a$ which allows the recursive definition for each $t>a$
$F_{t}(n)$ $=\frac{l+l(n-1)+t}{l(n-1)+t}F_{t}(n-1)$
$=\underbrace{\frac{ln+t}{l(n-1)+t}\frac{l(n-1)+t}{l(n-2)+t}}_{=\frac{ln+t}{l(n-2)+t}}F_{t}(n-2)$
$=\cdots=\frac{ln+t}{t}F_{t}(0)$
$=(t+ln)\underbrace{\frac{1}{t}F_{t}(0)}_{=:c(t)}$
@@
that since $t\mapsto F_{t}(n)$ is continuous that $c(t)$ is continuous. We extend $c$ to a function on all of $\mathbb{R}$ by using its periodicity.
@@
$\frac{F_{t+h}(0)-F_{t-h}(0)}{2h}$ $=\left(t-h\right)\frac{c(t+h)-c(t-h)}{2h}+c(t+h).$
Thus, by Lebesgue’s differentiation theorem, we then conclude that $f(t)=\mathbb{I}_{[0,a]}(t)g(t)+p(t)+tq(t)$ with $p,q$ $l$-periodic, and $p(t)=C+\int_{0}^{t}q(x)\textnormal{d}x$ for all $t>0$ and $g$ arbitrary. In particular, as $p$ is $l$-periodic we have $0=p(l)-p(0)=\int_{0}^{l}q(x)\textnormal{d}x$.
@@
As before we conclude that form $|f(t)|\leq 1$ that $q(t)=0$ and hence $p(t)=C$ is constant and $a^{-1}\int_{0}^{a}g(x)+p(x)\textnormal{d}x=l^{-1}\int_{a}^{a+l}p(x)\textnormal{d}x=C$ with $-C\leq g(t)\leq 1-C$. ∎